You need 1200mL of a 7% solution. You only stock a 12% concentration. How much of this 12% solution is needed?
Answer #1

Here we can imagine as

C1 = 7%

V1=1200ml

C2=12%

V2=?

C= concentration of fluid

V= volume of fluid

C1×V1 =C2×V2 here we want to know V2.

V2= C1×V1÷C2

If we apply the values

V2=7%×1200ml÷12%

V2=700ml.

Amswer is 700 ml